Show that the series , ∞Σ n=1 (x/1+n^2x^2) is uniformly convergent in [ k,1] for any k>0.

1
Expert's answer
2022-01-22T15:29:02-0500

If "x\in[k,1]" , "0<k<1" , then "\frac{x}{1+n^2x^2}\leq \frac{x}{n^2x^2}=\frac{1}{n^2x}\leq \frac{1}{n^2 k}".

The expression on the right hand does not depend on x and the series "\sum\limits_{n=1}^{\infty}\frac{1}{n^2 k}" is convergent, therefore the initial series is uniformly convergent on the segment [k,1].